OpenStudy (anonymous):

Which of the following is the factored form of the given equation? 6x^2 + 2x - 8 = 0

OpenStudy (anonymous):

2(3x + 4)(x - 1) = 0 (6x + 1)(x - 8) = 0 2(3x + 4)(x + 1) = 0

OpenStudy (anonymous):

= 2(3x^2+x-4) =2(3x+4)(x-1)

OpenStudy (anonymous):

ok explain please

OpenStudy (anonymous):

ok so its the first one

OpenStudy (anonymous):

you can find it by expand each of the equation..then you will get only the 1st equation can form the given equation. 2(3x + 4)(x - 1) = 0 (6x+8)(2x-2)=0 12x^2 +4x - 16 = 0 2(6x^2 + 2x -8) = 0

OpenStudy (anonymous):

then you can expand the 2nd and 3rd, but you will not get the same as the equation given.

OpenStudy (anonymous):

so, the answer is 2(3x + 4)(x - 1) = 0

OpenStudy (radar):

The 2 need only be expanded in the (3x+4) putting it in both terms result in it being expanded twice.

OpenStudy (anonymous):

ok thanks guys

OpenStudy (radar):

It would be much clear if you expanded the second two terms getting 2(3x^2+x-4) then expand the with the 2 getting 6x^2 + 2x -8